OpenStudy (fairytail_lover):

plzz helpp A biker traveled 24 miles in 3 hours. A unit rate to describe this would be 8 miles per hour. What is another unit rate in this situation? 1 over 8 mile per hour 1 over 8 hour per mile 1 mile per hour 3 hours per mile

OpenStudy (asylum15):

I'll solve this for you fairytail, just a second, bit rusty!

OpenStudy (fairytail_lover):

thank yew so much

OpenStudy (asylum15):

\[\frac{ 24 miles }{ 3 hours } = 8mph\] Invert it: \[\frac{ 3hours }{ 24miles } = \frac{ 3 }{ 24 } = \frac{ 1 }{ 8 } \] hours per mile.

OpenStudy (fairytail_lover):

omg thanksss !!